OpenStudy (anonymous):

How do I arrange the following physics equation to get R as slope and what do I graph (x and y)? I=sqrt(W/Rt)

OpenStudy (turingtest):

\[I=\sqrt{\frac{\omega}{Rt}}\]?

OpenStudy (anonymous):

yea

OpenStudy (turingtest):

what do you have after squaring both sides?

OpenStudy (anonymous):

I^2=W/Rt

OpenStudy (turingtest):

so you see that I is not linearly proportional to R hence, what do you mean you want R as a "slope"?

OpenStudy (anonymous):

We are required to graph this equation and find R (resistance) with the slope of the graph. I have values, and I can do that only after I solve this equation to get it into y=mx+b form.

OpenStudy (experimentx):

lol ... graph I^2 vs 1/R

OpenStudy (anonymous):

I can't. I'm trying to find R. I have the values for the other variables. I don't have R, so I can't graph it.

OpenStudy (experimentx):

what are you variables?

OpenStudy (anonymous):

I (current), W (energy), t (time)

OpenStudy (anonymous):

The equation is listed in the question.

OpenStudy (experimentx):

you can have only two variables independent and dependent on independent.

OpenStudy (anonymous):

time is constant. 300s. I and W have certain values

OpenStudy (experimentx):

what you are changing? what you are measuring keeping what constant?

OpenStudy (anonymous):

experimentX, I need to arrange the equation listed in y=mx+b form with R being the slope. Values are of no use right now.

OpenStudy (experimentx):